T(x1,x2,x3) = (x1+x2, x2+x3, x1-x3, 2x1+x2-x3) find he dimension of the kernel


1
Expert's answer
2021-01-12T16:06:46-0500

Let us find the kernel:


"\ker (T)=\{(x_1,x_2,x_3)\in\mathbb R^3\ |\ T(x_1,x_2,x_3)=(0,0,0,0)\}="

"=\{(x_1,x_2,x_3)\in\mathbb R^3\ |\ (x_1+x_2, x_2+x_3, x_1-x_3, 2x_1+x_2-x_3)=(0,0,0,0)\}"


It follows that we must solve the following system:


"\begin{cases}
x_1+x_2=0\\ x_2+x_3=0\\ x_1-x_3=0\\ 2x_1+x_2-x_3=0
\end{cases}" which is equivalent to the system


"\begin{cases}
x_1=-x_2\\ x_2=-x_3\\ x_1=x_3\\ 2x_3-x_3-x_3=0
\end{cases}" and to the system


"\begin{cases}
 x_2=-x_3\\ x_1=x_3\\x_3\in\mathbb R
\end{cases}"


Therefore, "\ker (T)=\{(x_1,x_2,x_3)\in\mathbb R^3\ |\ x_2=-x_3,x_1=x_3\}=\{(x_3,-x_3,x_3)\ |\ x_3\in\mathbb R\}=\{x_3(1,-1,1)\ |\ x_3\in\mathbb R\}."

It follows that "\dim(\ker(T))=1."